# KernJC: Automated Vulnerable Environment Generation for Linux Kernel Vulnerabilities

### v2

@@ 5.3. Role of Identified Configs
As shown in Figure 8, the vulnerability configs identified by KernJC are categorized into six categories: configs in vulnerability descriptions (), path-level configs (), code-level configs (), configs recursively reachable from direct configs (), configs with a one-hop select relationship to any direct config (), and configs with a one-hop depend relationship to any direct config (). KernJC can efficiently identify , , , , and through patch parsing and graph analysis. Comparatively, as demonstrated in § 2.2, the discovery of and is more challenging but essential to activating vulnerabilities in complex subsystems.
@@ 5.5.1. Vulnerabilities in BPF & Netfilter Subsystems
For example, although the basic CONFIG_BPF () necessitated by the six eBPF vulnerabilities (CVE-2016-4557, CVE-2017-16995, CVE-2020-27194, CVE-2020-8835, CVE-2021-34866, and CVE-2021-3490), as indicated by file dependencies in the Makefile, is easy to figure out with manual analysis, the examination of the six reveals a consistent requirement for the CONFIG_BPF_SYSCALL ( identified by KernJC) to be enabled for their activation, which is comparatively non-intuitive. The absence of CONFIG_BPF_SYSCALL will lead to wasted time and failed reproduction.
The analysis of the six Netfilter vulnerabilities reveals varied config requirements for each, along with their associated PoCs, complicating the manual construction of reproduction environments. Utilizing KernJC’s identification results, we conducted a manual analysis to determine the minimal config set of and for each vulnerability. Our findings indicate that the effective hidden configs for these vulnerabilities are exclusively derived from their respective sets, as detailed in Table 5. Specifically, we observed that the official announcement for CVE-2022-32250 (ANGUELKOV, [n. d.]) only cited the CONFIG_NETFILTER and CONFIG_NF_TABLES configs as necessary for activating this vulnerability. However, our reproduction demonstrated the additional requirement of CONFIG_NF_TABLES_IPV4. Similarly, the disclosure for CVE-2023-32233 (Krysiuk, [n. d.]) did not include the essential CONFIG_NFT_LOG and CONFIG_NFT_QUOTA config. The absence of these configs, as identified by KernJC, results in the failure of the original PoCs provided in these announcements during our reproduction experiments.
@@ 6.1. Reproducibility and Exploitability
Reproducibility and exploitability serve as vital metrics for gauging the severity of vulnerabilities and the associated risk in potentially susceptible environments. KernJC’s methodology aids in the reproducibility assessment of Linux kernel vulnerabilities by creating an environment where the specific vulnerability is both present and accessible from user space. However, the environments generated by KernJC may not fulfill the prerequisites for successful exploitation. For instance, to reliably exploit vulnerabilities in a modern Linux kernel, where various mitigation techniques are in place, many exploitations utilize the Filesystem in Userspace (FUSE) (LukeGix, [n. d.]) or userfaultfd (Dehors, [n. d.]). These mechanisms are contingent on the activation of non-default configs: CONFIG_FUSE_FS for FUSE and CONFIG_USERFAULTFD for userfaultfd. Since these configs are not related to the activation of vulnerabilities, their identification falls beyond the scope of this paper, which we intend to explore in future work.

### v3

@@ 5.3. Role of Identified Configs
As shown in Figure 8, the vulnerability configs identified by KernJC are categorized into six categories: configs in vulnerability descriptions (), path-level configs (), code-level configs (), configs recursively reachable from direct configs (), configs with a one-hop select
relationship to any direct config (), and configs with a one-hop depend
relationship to any direct config (). KernJC can efficiently identify , , , , and through patch parsing and graph analysis. Comparatively, as demonstrated in § 2.2, the discovery of and is more challenging but essential to activating vulnerabilities in complex subsystems.
@@ 5.5.1. Vulnerabilities in BPF & Netfilter Subsystems
For example, although the basic CONFIG_BPF
() necessitated by the six eBPF vulnerabilities (CVE-2016-4557, CVE-2017-16995, CVE-2020-27194, CVE-2020-8835, CVE-2021-34866, and CVE-2021-3490), as indicated by file dependencies in the Makefile, is easy to figure out with manual analysis, the examination of the six reveals a consistent requirement for the CONFIG_BPF_SYSCALL
( identified by KernJC) to be enabled for their activation, which is comparatively non-intuitive. The absence of CONFIG_BPF_SYSCALL
will lead to wasted time and failed reproduction.
The analysis of the six Netfilter vulnerabilities reveals varied config requirements for each, along with their associated PoCs, complicating the manual construction of reproduction environments. Utilizing KernJC’s identification results, we conducted a manual analysis to determine the minimal config set of and for each vulnerability. Our findings indicate that the effective hidden configs for these vulnerabilities are exclusively derived from their respective sets, as detailed in Table 5. Specifically, we observed that the official announcement for CVE-2022-32250 (ANGUELKOV, [n. d.]) only cited the CONFIG_NETFILTER
and CONFIG_NF_TABLES
configs as necessary for activating this vulnerability. However, our reproduction demonstrated the additional requirement of CONFIG_NF_TABLES_IPV4
. Similarly, the disclosure for CVE-2023-32233 (Krysiuk, [n. d.]) did not include the essential CONFIG_NFT_LOG
and CONFIG_NFT_QUOTA
config. The absence of these configs, as identified by KernJC, results in the failure of the original PoCs provided in these announcements during our reproduction experiments.
@@ 6.1. Reproducibility and Exploitability
Reproducibility and exploitability serve as vital metrics for gauging the severity of vulnerabilities and the associated risk in potentially susceptible environments. KernJC’s methodology aids in the reproducibility assessment of Linux kernel vulnerabilities by creating an environment where the specific vulnerability is both present and accessible from user space. However, the environments generated by KernJC may not fulfill the prerequisites for successful exploitation. For instance, to reliably exploit vulnerabilities in a modern Linux kernel, where various mitigation techniques are in place, many exploitations utilize the Filesystem in Userspace (FUSE) (LukeGix, [n. d.]) or userfaultfd (Dehors, [n. d.]). These mechanisms are contingent on the activation of non-default configs: CONFIG_FUSE_FS
for FUSE and CONFIG_USERFAULTFD
for userfaultfd. Since these configs are not related to the activation of vulnerabilities, their identification falls beyond the scope of this paper, which we intend to explore in future work.